What is the equation of through (-1,-5) slope =1
Answer by ikleyn(52781)

.

There are several (more than one) forms of such equations.


One of the forms is

    y-(-5) = 1*(x-(-1)),

or

    y + 5 = x + 1.

Solved.

------------------

At this site, there is a lesson
    - Equation for a straight line having a given slope and passing through a given point
where you can find many similar solved problems with detailed explanations to learn from.

Read it and learn the entire subject once and for all.